# Commonsense Reasoning in Arab Culture

### v1

@@ Abstract
Abdelrahman Sadallah1  Junior Cedric Tonga1  Khalid Almubarak2 Saeed Almheiri1Farah Atif1Cahtrine Qwaider1Karima Kadaoui1 Sara Shatnawi3Yaser Alesh4Fajri Koto1 1Department of Natural Language Processing, MBZUAI 2SDAIA   3Al-Balqa Applied University   4Khalifa University {abdelrahman.sadallah,fajri.koto}@mbzuai.ac.ae
@@
Abdelrahman Sadallah1  Junior Cedric Tonga1  Khalid Almubarak2 Saeed Almheiri1Farah Atif1Cahtrine Qwaider1Karima Kadaoui1 Sara Shatnawi3Yaser Alesh4Fajri Koto1 1Department of Natural Language Processing, MBZUAI 2SDAIA   3Al-Balqa Applied University   4Khalifa University {abdelrahman.sadallah,fajri.koto}@mbzuai.ac.ae
@@ 1 Introduction
Figure 1: ArabCulture covers four regions across the Middle East and North Africa, spanning 13 countries. The highlighted areas on the map represent the regions included in ArabCulture. We present example questions from Saudi Arabia, Jordan, Egypt, and Morocco, each with three answer choices, with the correct answer marked in bold. English translations are provided for illustration.
@@
The Arab world, home to approximately 456 million people (Diab et al., 2017; Shoufan and Alameri, 2015), is characterized by its linguistic unity through Modern Standard Arabic (MSA) while encompassing diverse traditions, religions, and customs (Mirkin, 2010). This cultural diversity influences not only social interactions but also reasoning patterns, making it crucial to develop evaluation benchmarks that reflect these variations. However, most existing commonsense reasoning datasets are developed with Western-centric assumptions, limiting their applicability to Arabic-speaking societies. Despite recent advancements in Arabic LLMs (Sengupta et al., 2023; Huang et al., 2023; Team, 2024; Bari et al., 2024b), their evaluation has largely relied on machine-translated datasets originally created in English. Many commonsense reasoning benchmarks (Bisk et al., 2019; Sap et al., 2019a) fail to capture Arab cultural perspectives, as simple translations do not account for region-specific knowledge, potentially introducing bias. Given the significant cultural variations across the Arab world, these datasets do not provide a holistic measure of Arabic LLMs’ ability to reason within culturally specific contexts. This raises an important question: To what extent can existing LLMs accurately reason about commonsense knowledge in diverse cultural settings, particularly in the Arab world? To address this gap, we introduce ArabCulture, a commonsense reasoning dataset specifically designed to assess the cultural knowledge of Arabic LLMs. The dataset consists of 3,482 questions written in Modern Standard Arabic (MSA), covering 13 countries across the Gulf, Levant, North Africa, and the Nile Valley (see Figure 1). It spans 12 major domains and 54 fine-grained subtopics, reflecting various aspects of social norms, traditions, and daily life in the Arab world. Unlike existing benchmarks, which often rely on translated datasets, ArabCulture was built from scratch by directly engaging native speakers to write culturally relevant questions for their respective countries. We carefully implemented quality control measures throughout the dataset creation process, including rigorous validation steps to maintain accuracy, relevance, and cultural sensitivity. We evaluate a range of closed-weight and open-weight Arabic and multilingual LLMs in a zero-shot setting to assess their cultural commonsense reasoning capabilities. Inspired by Koto et al. (2024b), we frame the task in two ways: multiple-choice questions (MCQ) and completion tasks. Additionally, we introduce three levels of location-based contextual grounding: (1) no additional location information, (2) specifying the broader region (e.g., Gulf or Levant), and (3) specifying the exact country along with its regional classification. This setup allows us to analyze how effectively LLMs incorporate geographical and cultural cues in their reasoning. Our results show that even LLMs with up to 32B parameters struggle with cultural commonsense reasoning, with performance varying significantly across regions. We conduct a detailed analysis of the best-performing models, identifying strengths and weaknesses across different cultural contexts. We also conducted a small manual experiment to test the models’ ability to explain their chosen answers. Additionally, we explore whether enriching prompts with cultural facts improves performance in smaller language models, finding that while it helps in some cases, it does not provide a universal solution.
@@ 2.1 Commonsense Reasoning in English
Early research on commonsense reasoning primarily focused on linguistic reasoning, as seen in the Winograd Schema Challenge (Levesque et al., 2012) and Winogrande (Sakaguchi et al., 2019), which evaluate pronoun coreference resolution within social and linguistic contexts. Other works have explored physical commonsense reasoning, assessing model’s understanding of real-world properties and relationships (Bisk et al., 2019), as well as social reasoning, where models are tested on their ability to interpret human emotions, actions, and social norms (Sap et al., 2019b). Research has also expanded into numerical Lin et al. (2020); Akhtar et al. (2023), temporal Tan et al. (2023), and causal reasoning Roemmele et al. (2011); Du et al. (2022), broadening the scope of commonsense evaluation. However, these benchmarks are all developed in English and shaped by Western cultural assumptions, limiting their applicability to the Arab world.
@@ 2.2 Arabic Large Language Models and Their Evaluation on Commonsense Reasoning
A limited number of Arabic language models have been developed with more than 7B parameters, all of which are decoder-only architectures. These include JAIS (Sengupta et al., 2023), Fanar (Fanar-Team, 2024), AceGPT (Huang et al., 2023), and ALLAM (Bari et al., 2024b). Their evaluation of commonsense reasoning has been primarily based on machine-translated datasets from English to Arabic (e.g.,  Tawalbeh and Al-Smadi (2020); Al-Bashabsheh et al. (2021)). Although this approach provides a useful reference, it does not offer a comprehensive assessment of how well these models capture culturally grounded commonsense knowledge. Much of the recent Arabic-centric benchmarks have focused on classic NLP tasks, including syntax, semantics, and question answering, as seen in LaraBench (Abdelali et al., 2024), natural language generation in DOLPHIN (Elmadany et al., 2023a), and natural language understanding in ORCA (Elmadany et al., 2023b). Only a few studies have shifted their focus toward knowledge-intensive tasks and reasoning abilities. Among them, ArabicMMLU (Koto et al., 2024a) compiles exam questions from different education levels across Arabic-speaking countries, offering a broad knowledge assessment but placing less emphasis on cultural reasoning. Table 1 compares ArabCulture with related Arabic datasets. Most existing datasets are derived from machine translation with post-editing, prioritizing linguistic accuracy over cultural relevance. While some assess reasoning, they often lack cultural grounding, location metadata, and fine-grained topic categorization. ACVA (Huang et al., 2023), generated using ChatGPT (Ouyang et al., 2022), is not designed for reasoning evaluation. Similarly, AraDice-Culture (Mousi et al., 2025) consists of only 180 samples and focuses on open-ended cultural knowledge rather than structured reasoning tasks. These gaps highlight the need for larger, more diverse benchmarks that better capture Arabic cultural contexts and reasoning abilities.
@@ 4.1 Experimental Setup
We conducted zero-shot experiments across 31 models, categorized into the following groups: (1) 20 multilingual models of various sizes, such as BLOOMZ Muennighoff et al. (2023), mT0-xxl Muennighoff et al. (2023), Llama–3 Grattafiori et al. (2024), Aya-Expanse Dang et al. (2024), Gemma-2 Riviere et al. (2024), and Qwen2.5 Qwen et al. (2025); (2) 10 Arabic-centric models of different sizes, including Jais Sengupta et al. (2023), SILMA Team (2024), AceGPT-v2 Liang et al. (2024), and ALLaM Bari et al. (2024a); (3) 1 closed-weight model, GPT-4o OpenAI et al. (2024a). All experiments are done using zero temperature (greedy sampling) to enforce the models to produce factual outputs. We conducted experiments using both Arabic and English prompts ( Figure 5) and evaluated language models using two strategies: (1) sentence completion and (2) MCQ. In the sentence completion approach, we concatenate the premise with each candidate’s second sentence and select the one with the highest likelihood. For MCQ, we assign alphabetical labels to the answer choices (A, B, C for English, and \<أ> , \<ب>, \<ج> for Arabic), and the selected answer corresponds to the option with the highest probability. We constructed these experiments using the LM-Evaluation-Harness Framework Gao et al. (2024). Note that for the closed-weight model, we only perform MCQ-style evaluation, instructing the model to generate the answer as a JSON object containing only the answer character. As discussed in Section 1, cultural knowledge varies across locations, and we hypothesize that providing geographical context can enhance a model’s reasoning ability in cultural contexts. To test this, we complement our experiments with three different levels of location context {none,region,region + $\ell\in\{\text{none},\text{region},\text{region $+$ country}\}$.
@@ 4.2 Zero-shot Experiments
The overall results presented in Table 3 reveal notable performance differences between open-weight and closed-weight models in understanding Arabic culture and norms. While some large-scale open-weight models achieve relatively high accuracy—such as Qwen-2.5-Instruct (72B) with 80%, LLaMA-3.3-Instruct (70B) with 75.4%, and AceGPT-v2-Chat (32B) with 79.7%—, closed-wight models represented by GPT-4o demonstrate significantly stronger performance. Arabic-centric models do not consistently outperform multilingual models. Some, such as Jais, struggle despite being tailored for Arabic, whereas certain multilingual models like Qwen2.5 and LLama-3.3 Instruct surpass them in accuracy. Within the same model family, performance generally improves with larger model sizes—except for Jais and AceGPT. However, across different model families, scaling does not guarantee better results. This indicates that factors beyond model size, such as pretraining data, architecture, and training recipes, significantly impact cultural comprehension. When comparing base models to instruction-tuned variants, we observe modest improvements in completion tasks but substantial gains in MCQ tasks. Overall, these findings highlight the need for improvements in Arabic and multilingual models to enhance their comprehension of Arabic cultural contexts. Addressing this gap can also help mitigate cultural biases, which have been identified in recent studies, such as the work by Naous et al. (2024).
@@ C.2 Cultural Context prompts
Figure 6 shows the Arabic that we used to generate the five cultural sentences for improving the small language model with additional context from GPT-4o. A translation in English is available to allow those who do not understand Arabic to comprehend the prompt.
@@ Appendix D Results of the Arabic prompt in the zero-shot experiments
Table 9 includes the zero-shot experiment results for the Arabic Prompt. The prompt is shown in Figure 5.
@@ D.1 Results by Geographic Location - Arabic Prompt
Table 6 includes the geographical analysis results for the Arabic Prompt.
@@ D.2 Results by Topic - Arabic Prompt
Table 7 includes the analysis by topic results for the Arabic Prompt.

### v2

@@ Abstract
Abdelrahman Sadallah1  Junior Cedric Tonga1  Khalid Almubarak2,5 Saeed Almheiri1Farah Atif1Chatrine Qwaider1Karima Kadaoui1 Sara Shatnawi3Yaser Alesh4Fajri Koto1 1Mohamed bin Zayed University of Artificial Intelligence 2SDAIA   3Al-Balqa Applied University 4Khalifa University   5HUMAIN, Data and AI Models {abdelrahman.sadallah,fajri.koto}@mbzuai.ac.ae
@@
Abdelrahman Sadallah1  Junior Cedric Tonga1  Khalid Almubarak2,5 Saeed Almheiri1Farah Atif1Chatrine Qwaider1Karima Kadaoui1 Sara Shatnawi3Yaser Alesh4Fajri Koto1 1Mohamed bin Zayed University of Artificial Intelligence 2SDAIA   3Al-Balqa Applied University 4Khalifa University   5HUMAIN, Data and AI Models {abdelrahman.sadallah,fajri.koto}@mbzuai.ac.ae
@@ 1 Introduction
Figure 1: ArabCulture covers four regions across the Middle East and North Africa, spanning 13 countries. The highlighted areas on the map represent the regions included in ArabCulture. We present example questions for the Lunch category from Algeria, Egypt, Syria, and KSA, each with three answer choices, with the correct answer marked in bold. English translations are provided for illustration.
@@
The Arab world, home to approximately 456 million people (Diab et al., 2017; Shoufan and Alameri, 2015), is characterized by its linguistic unity through Modern Standard Arabic (MSA) while encompassing diverse traditions, religions, and customs (Mirkin, 2010). This cultural diversity influences not only social interactions but also reasoning patterns, making it crucial to develop evaluation benchmarks that reflect these variations. However, most existing commonsense reasoning datasets are developed with Western-centric assumptions, limiting their applicability to Arabic-speaking societies. Despite recent advancements in Arabic LLMs (Sengupta et al., 2023; Huang et al., 2024; Team, 2024; Bari et al., 2024b), their evaluation has largely relied on machine-translated datasets originally created in English. Many commonsense reasoning benchmarks (Bisk et al., 2019; Sap et al., 2019a) fail to capture Arab cultural perspectives, as simple translations do not account for region-specific knowledge, potentially introducing bias. Given the significant cultural variations across the Arab world, these datasets do not provide a holistic measure of Arabic LLMs’ ability to reason within culturally specific contexts. This raises an important question: To what extent can existing LLMs accurately reason about commonsense knowledge in diverse cultural settings, particularly in the Arab world? To address this gap, we introduce ArabCulture, a commonsense reasoning dataset specifically designed to assess the cultural knowledge of Arabic LLMs. The dataset consists of 3,482 questions written in Modern Standard Arabic (MSA), covering 13 countries across the Gulf, Levant, North Africa, and the Nile Valley (see Figure 1). It spans 12 major domains and 54 fine-grained subtopics, reflecting various aspects of social norms, traditions, and daily life in the Arab world. Unlike existing benchmarks, which often rely on translated datasets, ArabCulture was built from scratch by directly engaging native speakers to write culturally relevant questions for their respective countries. We carefully implemented quality control measures throughout the dataset creation process, including rigorous validation steps to maintain accuracy, relevance, and cultural sensitivity. We evaluate a range of closed-weight and open-weight Arabic and multilingual LLMs in a zero-shot setting to assess their cultural commonsense reasoning capabilities. Inspired by Koto et al. (2024b), we frame the task in two ways: multiple-choice questions (MCQ) and completion tasks. Additionally, we introduce three levels of location-based contextual grounding: (1) no additional location information, (2) specifying the broader region (e.g., Gulf or Levant), and (3) specifying the exact country along with its regional classification. This setup allows us to analyze how effectively LLMs incorporate geographical and cultural cues in their reasoning. Our results show that even LLMs with up to 32B parameters struggle with cultural commonsense reasoning, with performance varying significantly across regions. We conduct a detailed analysis of the best-performing models, identifying strengths and weaknesses across different cultural contexts. We also conducted a small manual experiment to test the models’ ability to explain their chosen answers. Additionally, we explore whether enriching prompts with cultural facts improves performance in smaller language models, finding that while it helps in some cases, it does not provide a universal solution.
@@ 2.1 Commonsense Reasoning in English
Early research on commonsense reasoning primarily focused on linguistic reasoning, as seen in the Winograd Schema Challenge (Levesque et al., 2012) and Winogrande (Sakaguchi et al., 2021), which evaluate pronoun coreference resolution within social and linguistic contexts. Other works have explored physical commonsense reasoning, assessing model’s understanding of real-world properties and relationships (Bisk et al., 2019), as well as social reasoning, where models are tested on their ability to interpret human emotions, actions, and social norms (Sap et al., 2019b). Research has also expanded into numerical Lin et al. (2020); Akhtar et al. (2023), temporal Tan et al. (2023), and causal reasoning Roemmele et al. (2011); Du et al. (2022), broadening the scope of commonsense evaluation. However, these benchmarks are all developed in English and shaped by Western cultural assumptions, limiting their applicability to the Arab world.
@@ 2.2 Arabic Large Language Models and Their Evaluation on Commonsense Reasoning
A limited number of Arabic language models have been developed with more than 7B parameters, all of which are decoder-only architectures. These include JAIS (Sengupta et al., 2023), Fanar (Fanar-Team, 2024), AceGPT (Huang et al., 2024), and ALLAM (Bari et al., 2024b). Their evaluation of commonsense reasoning has been primarily based on machine-translated datasets from English to Arabic (e.g.,  Tawalbeh and Al-Smadi (2020); Al-Bashabsheh et al. (2021)). Although this approach provides a useful reference, it does not offer a comprehensive assessment of how well these models capture culturally grounded commonsense knowledge. Much of the recent Arabic-centric benchmarks have focused on classic NLP tasks, including syntax, semantics, and question answering, as seen in LaraBench (Abdelali et al., 2024), natural language generation in DOLPHIN (Elmadany et al., 2023a), and natural language understanding in ORCA (Elmadany et al., 2023b). Only a few studies have shifted their focus toward knowledge-intensive tasks and reasoning abilities. Among them, ArabicMMLU (Koto et al., 2024a) compiles exam questions from different education levels across Arabic-speaking countries, offering a broad knowledge assessment but placing less emphasis on cultural reasoning. Table 1 compares ArabCulture with related Arabic datasets. Most existing datasets are derived from machine translation with post-editing, prioritizing linguistic accuracy over cultural relevance. While some assess reasoning, they often lack cultural grounding, location metadata, and fine-grained topic categorization. ACVA (Huang et al., 2024), generated using ChatGPT (Ouyang et al., 2022), is not designed for reasoning evaluation. Similarly, AraDice-Culture (Mousi et al., 2025) consists of only 180 samples and focuses on open-ended cultural knowledge rather than structured reasoning tasks. These gaps highlight the need for larger, more diverse benchmarks that better capture Arabic cultural contexts and reasoning abilities.
@@ 3.2 Data Statistics
DeepSeek-R1-Distill-Llama (70B) 37.4 37.6 38.4 34.3 35.2 34.5
DeepSeek-R1-Distill-Qwen (32B) 34.8 34.7 35.0 34.3 34.3 34.3
QwQ (32B) 34.4 35.7 36.4 36.7 35.1 35.6
@@ 4.1 Experimental Setup
We conducted zero-shot experiments across 31 models, categorized into the following groups: (1) 20 multilingual models of various sizes, such as BLOOMZ Muennighoff et al. (2023), mT0-xxl Muennighoff et al. (2023), Llama–3 Grattafiori et al. (2024), Aya-Expanse Dang et al. (2024), Gemma-2 Riviere et al. (2024), and Qwen2.5 Qwen et al. (2025); (2) 10 Arabic-centric models of different sizes, including Jais Sengupta et al. (2023), SILMA Team (2024), AceGPT-v2 Huang et al. (2024), and ALLaM Bari et al. (2024a); (3) 3 reasoning models that are distilled from the DeepSeek-R1 model (DeepSeek-AI et al., 2025). (4) 1 closed-weight model, GPT-4o OpenAI et al. (2024a). All experiments are done using zero temperature (greedy sampling) to enforce the models to produce factual outputs. We conducted experiments using both Arabic and English prompts ( Figure 5) and evaluated language models using two strategies: (1) sentence completion and (2) MCQ. In the sentence completion approach, we concatenate the premise with each candidate’s second sentence and select the one with the highest likelihood. For MCQ, we assign alphabetical labels to the answer choices (A, B, C for English, and \<أ> , \<ب>, \<ج> for Arabic), and the selected answer corresponds to the option with the highest probability. We constructed these experiments using the LM-Evaluation-Harness Framework Gao et al. (2024). Note that for the closed-weight model, we only perform MCQ-style evaluation, instructing the model to generate the answer as a JSON object containing only the answer character. As discussed in Section 1, cultural knowledge varies across locations, and we hypothesize that providing geographical context can enhance a model’s reasoning ability in cultural contexts. To test this, we complement our experiments with three different levels of location context {none,region,region + $\ell\in\{\text{none},\text{region},\text{region $+$ country}\}$.
@@ 4.2 Zero-shot Experiments
The overall results presented in Table 3 reveal notable performance differences between open-weight and closed-weight models in understanding Arabic culture and norms. While some large-scale open-weight models achieve relatively high accuracy—such as Qwen-2.5-Instruct (72B) with 80%, LLaMA-3.3-Instruct (70B) with 75.4%, and AceGPT-v2-Chat (32B) with 79.7%—, closed-wight models represented by GPT-4o demonstrate significantly stronger performance. Arabic-centric models do not consistently outperform multilingual models. Some, such as Jais, struggle despite being tailored for Arabic, whereas certain multilingual models like Qwen2.5 and Llama-3.3 Instruct surpass them in accuracy. Within the same model family, performance generally improves with larger model sizes—except for Jais and AceGPT. However, across different model families, scaling does not guarantee better results. This indicates that factors beyond model size, such as pretraining data, architecture, and training recipes, significantly impact cultural comprehension. When comparing base models to instruction-tuned variants, we observe modest improvements in completion tasks but substantial gains in MCQ tasks. While reasoning models have demonstrated impressive capabilities in domains such as mathematics and programming, they perform poorly on our cultural reasoning tasks. This discrepancy suggests that cultural reasoning involves fundamentally different challenges that remain underexplored and require dedicated attention. Overall, these findings highlight the need for improvements in Arabic and multilingual models to enhance their comprehension of Arabic cultural contexts. Addressing this gap can also help mitigate cultural biases, which have been identified in recent studies, such as the work by Naous et al. (2024).
@@ 6 Conclusion
Acknowledgments
We are grateful to the supercomputing center at MBZUAI for providing resources for this research. We also gratefully acknowledge the support of the university for providing funding to hire the human annotators.
@@ C.2 Cultural Context prompts
Figure 6 displays the Arabic prompt used to generate five culturally grounded sentences for enhancing the small language model with additional context from GPT-4o. An English translation is provided to assist readers unfamiliar with Arabic in understanding the prompt.
@@ Appendix D Results of the Arabic prompt in the zero-shot experiments
Table 9 presents the zero-shot experiment results using the Arabic prompt, which is illustrated in Figure 5.
@@ D.1 Results by Geographic Location - Arabic Prompt
Table 6 presents the country-level breakdown analysis based on the Arabic prompt.
@@ D.2 Results by Topic - Arabic Prompt
Table 7 presents the breakdown analysis by topic based on the Arabic prompt.
@@
DeepSeek-R1-Distill-Llama (70B) 36.7 37.1 37.6 34.7 34.5 35.0
DeepSeek-R1-Distill-Qwen (32B) 33.6 34.4 35.0 34.3 34.3 34.3
QwQ (32B) 22.83 22.29 22.77 32.88 32.62 32.28